At 273K and 10 bar, the Langmuir adsorption of a gas on a solid surface gave the fraction of surface coverage as 0.01. The Langmuir adsorption isotherm constant is ______

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Langmuir adsorption gives the formulae:
L = KP/(1+KP)
Putting pressure and L values we get,
0.01 = 10K/(1+10K)
 
0.01 + 0.1K = 10K
9.9K = 0.01
K = 0.01/9.9 => 1/990